OpenStudy (anonymous):

The tax on a meal costing $16 is $1.20. At the same tax rate, what is the tax on a meal costing $22? Round your answer to the nearest cent.

jimthompson5910 (jim_thompson5910):

solve 1.2/16 = x/22 for x

OpenStudy (anonymous):

7.5

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

1.2/16 = x/22 1.2*22 = 16x x = ???

OpenStudy (anonymous):

1.75???

jimthompson5910 (jim_thompson5910):

how did you get that

OpenStudy (anonymous):

idk im so confused

jimthompson5910 (jim_thompson5910):

how would you isolate x in 1.2*22 = 16x

OpenStudy (anonymous):

idk math im really dumb wen it cums to math ;(

jimthompson5910 (jim_thompson5910):

what undoes multiplication?

OpenStudy (anonymous):

1.65 ???????

jimthompson5910 (jim_thompson5910):

correct